OpenStudy (anonymous):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? Answer {-13, -3, 2, 3, 13} {12, 3, 0, 3, 13} {-17, -7, -2, 3, 13}

OpenStudy (anonymous):

The domain refers to the x values, while the range refers to the y values. So you'll find the range of y = 5x - 2 by substituting in each of the domain instead of x in the equation ex: y=5*-3-2=-17 and so on....... ___ U got it ?

OpenStudy (anonymous):

noo,

OpenStudy (anonymous):

You will substitute every number in the domain set in the equation to get the range :D Put instead of x the domains number y=5*-3-2=-17 y=5*-1-2=-7 y=5*0-2=-2 y=5*1-2=3 y=5*3-2=13 Therefore the range ={-17, -7, -2, 3, 13}.. Make sense ?

OpenStudy (anonymous):

yesssss thanks!

OpenStudy (anonymous):

Yw :)